# HumanVBench: Exploring Human-Centric Video Understanding Capabilities of MLLMs with Synthetic Benchmark Data

### v2

@@ Abstract
In the domain of Multimodal Large Language Models (MLLMs), achieving human-centric video understanding remains a formidable challenge. Existing benchmarks primarily emphasize object and action recognition, often neglecting the intricate nuances of human emotions, behaviors, and speech-visual alignment within video content. We present HumanVBench, an innovative benchmark meticulously crafted to bridge these gaps in the evaluation of video MLLMs. HumanVBench comprises 16 carefully designed tasks that explore two primary dimensions: inner emotion and outer manifestations, spanning static and dynamic, basic and complex, as well as single-modal and cross-modal aspects. With two advanced automated pipelines for video annotation and distractor-included QA generation, HumanVBench utilizes diverse state-of-the-art (SOTA) techniques to streamline benchmark data synthesis and quality assessment, minimizing human annotation dependency tailored to human-centric multimodal attributes. A comprehensive evaluation across 22 SOTA video MLLMs reveals notable limitations in current performance, especially in cross-modal and emotion perception, underscoring the necessity for further refinement toward achieving more human-like understanding. HumanVBench is open-sourced to facilitate future advancements and real-world applications in video MLLMs.
Figure 1: Overview of HumanVBench, which encompasses 16 fine-grained tasks for extensive human-centric evaluations (middle blue box). Each task is denoted by its acronym and the number of included QA instances. The right orange box illustrates some examples of these QAs. HumanVBench is constructed using the novel automated Video Annotation Pipeline (upper left, purple box), followed by the Distractor-Included QA Synthesis Pipeline (lower left, green box). These pipelines are reusable and backed by more than twenty data processing operators with advanced algorithm implementation and cutting-edge auxiliary models.
@@ 1 Introduction
In recent years, Multimodal Large Language Models (MLLMs) have emerged as pivotal technological advancements, significantly expanding traditional language model capabilities to include processing and comprehending diverse data forms like text, images, and videos [44, 8, 3]. Among these, video-oriented MLLMs [55, 11, 34] have garnered substantial research interest due to their potential in interpreting video content in a manner closely aligned with human perception. While image-based MLLMs [36, 38, 62, 50, 49, 24] primarily focus on static content, video-MLLMs offer enhanced capacities for understanding the complex temporal dynamics intrinsic to video data.
Human-centric scenes in videos naturally attract attention due to the emphasis on individuals’ emotions, actions, and verbal interactions, necessitating effective comprehension by video understanding models. Despite advances in this field, existing benchmarks often fall short in rigorously assessing the nuanced understanding of human emotions and behaviors. Current evaluations predominantly focus on general content comprehension, object recognition, and action detection, often neglecting subtle intricacies such as emotional insight and behavioral analysis. Furthermore, synchronizing speech with visual elements remains a substantial challenge; unlike humans, who effortlessly discern mismatches between audio and visual cues, computational models often struggle with tasks like identifying speakers and aligning speech with corresponding lip movements.
To bridge these gaps, we introduce HumanVBench, a pioneering benchmark specifically tailored for video MLLMs focusing on human-centric analysis. HumanVBench includes 16 meticulously designed QA tasks, categorized into two core aspects: inner emotion and outer manifestation, as illustrated in Figure 1. The inner emotion dimension evaluates the model’s capacity to perceive emotional cues from videos, while the outer manifestation encompasses person recognition, human behavior analysis, and cross-modal speech-visual alignment. Each category is further subdivided into many fine-grained tasks, offering a comprehensive evaluation across static and dynamic, basic and complex, single-modal and cross-modal aspects. For instance, it enables us to examine MLLMs’ understanding of human actions, emotions, and the associated causal relationships, as well as consider the alignment degree between appearance, mouth movement, audio, and speech content in videos.
The construction of HumanVBench is facilitated by the advanced open-source data processing system, Data-Juicer [6], and involves two novel data synthesis pipelines: the Human-Centric Video Annotation Pipeline and the Distractor-Included QA Synthesis Pipeline, leveraging over twenty state-of-the-art (SOTA) data processing operators. Unlike conventional benchmarks that rely heavily on human annotators [26, 39, 42, 51, 17, 18], our approach automates multi-modal, fine-grained annotation processes, with the help of cooperation by diverse task-specific annotation algorithms and models. The Distractor-Included QA Synthesis Pipeline uses iterative enhancement with video-MLLMs to refine question accuracy and generate distractors, while ensuring quality through a final human review. Thanks to the fruitful annotations and dedicated orchestrations of data processing operators, the construction of HumanVBench is largely automated, substantially reducing the need for manual intervention. Additionally, our method is applicable to “in-the-wild” video data, enabling the creation of video benchmarks that are not confined to controlled or domain-specific environments.
@@
Through HumanVBench, we comprehensively evaluate 22 SOTA video MLLMs, including open-source models like VideoLLaMA3 [57] and commercial ones like GPT-4o [41]. Our evaluations reveal several interesting insights and significant gaps between current model capabilities and human-like understanding, particularly in tasks such as cross-modal alignment and emotion perception. While proprietary models demonstrate closer human-like accuracy, open-source models frequently misclassify emotions due to temporal noise, underscoring the need for further architectural improvements and refined datasets.
@@
  • We introduce HumanVBench, a novel video benchmark for MLLMs that emphasizes fine-grained human comprehension in videos, focusing on emotion perception, person identification, behavioral analysis, and speech-visual alignment.
  • We propose two advanced pipelines for automatic video annotation and the production of high-quality, multiple-choice questions relevant to video-based descriptive queries. These pipelines reduce manual labor and are equipped with diverse reusable operators for multimodal, detailed individual labeling in videos, making them adaptable across various contexts.
  • Our comprehensive evaluation of numerous SOTA video MLLMs offers key insights, facilitating in-depth discussions regarding their performance, strengths, and areas for enhancement.
  • We release our benchmark, including data, evaluation, and synthesis codes at https://github.com/modelscope/data-juicer/tree/HumanVBench , to foster further evolution of future human-centric video analysis systems.
@@ 2 Related Works
Multimodal Large Language Models. The remarkable progress in Large Language Models (LLMs) has sparked extensive research into merging language comprehension with visual and auditory information, thereby expediting the advancement of multimodal models. Within this domain, image MLLMs amalgamate visual and linguistic data to enhance image interpretation and cross-modal reasoning [13, 32, 23, 37]. Video MLLMs [29, 39, 34, 48, 58, 16] extend these capabilities by incorporating temporal sequences for dynamic video analysis. Furthermore, generalist MLLMs [20, 19, 61] can process a plethora of inputs, including images, video, and audio, thus improving adaptive task performance across various modalities. Despite these advancements, rigorous evaluation of video MLLMs on human-centric video understanding tasks remains an unaddressed challenge, which this work aims to address.
Video Benchmarks for MLLMs. There exists a substantial array of benchmarks for assessing the performance of MLLMs, with a particular emphasis on video-based evaluations. Presently, general-purpose video benchmarks can be broadly categorized based on their evaluative methodologies: multiple-choice queries with definitive answers [40, 42, 30, 18, 45], Open-Ended Video Question Answering necessitating supplementary LLMs for evaluation [52, 39, 17, 45], video-captioning benchmarks [52, 12, 5], and video-generation assessments such as VBench [22]. Additionally, there are human-centric benchmarks in MLLMs. ActivityNet-QA [52], rooted in the widely-used ActivityNet dataset [56], emphasizes object and human activity recognition. HumanBench [46] amalgamates multiple public datasets to create benchmarks for tasks like pose estimation, pedestrian attribute recognition, and crowd assessment. HERM [31] introduces a human-centric image benchmark focusing on people, poses, actions, and interactions, using GPT-4-generated question-answer pairs.
Our Position. Compared to existing benchmarks, HumanVBench distinguishes itself through both innovative task design and unique construction methodologies. Firstly, it introduces novel human-related video tasks. Unlike HERM, HumanVBench integrates complex scenarios encompassing temporal dynamics for enhanced human video interpretation; compared to ActivityNet-QA and HumanBench, HumanVBench delves into nuanced emotional comprehension and evaluates cross-modal alignment between visual data and speech modalities. Moreover, HumanVBench boasts a pioneering construction approach, deriving from raw, uncurated video content and employing over 20 advanced processing operators for meticulous video character annotation and automatized multiple-choice question formulation. Human intervention is minimized to verify the quality of well-structured QAs.
@@ 3.1 Task Design and Definition
Human observers naturally concentrate on individuals in videos, examining their appearance, emotions, and behaviors. This intrinsic focus underpins our design of 16 fine-grained, human-centric tasks, aimed at evaluating MLLMs’ ability to mimic human-like perception and understanding in video analysis (Figure 1). Each task is detailed with definitions and examples in the Appendix Section 10. The tasks are grouped into two categories based on content observability: Inner Emotion and Outer Manifestation.
3.1.1 Inner Emotion
In real-world videos, inner emotions are less observable, as cameras often capture wider scenes instead of facial close-ups. Thus, emotion perception becomes an abstract skill requiring careful attention to facial expressions, body language, and verbal cues. With this goal, tasks in this category evaluate video-MLLMs’ ability to detect and interpret emotional nuances, mirroring human perceptual skills.
Specifically, five specific tasks comprise the Emotion Perception category:
  • Emotion Recognition (ER) requires identifying the most fitting emotional description for an individual in a video.
  • Emotion Temporal Analysis (ETA) focuses on tracking emotional changes over time.
  • Attitude Recognition (AT) assesses the inferred attitudes of individuals in relation to specific events or entities, classifying them as positive, negative, or neutral.
  • Emotion Intensity Comparison (EIC) evaluates the model’s ability to differentiate and quantify the emotional intensity of various individuals.
These tasks collectively enable the evaluation of MLLMs’ potential in capturing emotional cues and facial details from videos, thereby gauging their capabilities in inner emotional understanding.
@@ 3.1.1 Inner Emotion
Figure 3: The Distractor-Included QA Synthesis Pipeline facilitates four steps: selecting “question videos”, generating preliminary answers, optimizing answers with generated distractors, and manually verifying multiple-choice questions.
3.1.2 Outer Manifestation
Unlike inner emotions, outer manifestations deal with more tangible aspects such as identifying individual(s), causality reasoning, and synchronization of video elements like speech or singing. Guided by these aspects of human observation, we formulated three task categories:
The first basic dimension is for Person Recognition, evaluating the model’s capability to identify a particular person within a complex scene, akin to “person-finding”:
  • Text-to-Human (T2H) tests the model’s ability to identify a person based on textual descriptions.
  • Human-to-Text (H2T) assesses the accuracy of text explanations attributed to a target person, distinguishing them from others.
  • Human Counting (HC) evaluates the model’s ability to detect, track, and count distinct individuals in a video.
  • Appearance Time Detection (ATD) requires identifying a specified individual’s presence time in the video.
Besides, we incorporate the Human Behavior Analysis category to scrutinize the model’s ability to understand and analyze individual behaviors, with four tasks:
  • Behavior Temporal Analysis (BTA) explores behavior tracking capabilities over time.
  • Behavior Causality Analysis (BCA) examines causal inferences within behavior sequences.
  • Action at Specified Time (AST) requires identifying precise actions at given times.
  • Time of Specific Action (TSA) tests the detection of exact moments specifying actions occur.
Furthermore, a critical cross-modal category is Speech-Visual Alignment, designed to test capabilities involving the following tasks:
  • Audio-Visual Speaker Matching (AVSM) correlates audio features to identify individuals and analyzes matches by appearance (gender and age).
  • Active Speaker Detection (ASD) identifies the individual currently speaking by integrating visual and audio inputs.
  • Audio-Visual Alignment Detection (AVAD) detects synchronization points, examining the coherence between lip movements and audio.
  • Speech Content Matching (SCM) requires analyzing spoken content against text, and evaluating transcription or lip-reading capabilities.
These aforementioned tasks together enable a comprehensive assessment of MLLMs’ diverse perceptual abilities across emotion, identity, behavior and speech-visual alignment, pushing toward human-like video comprehension.
@@ 3.2 Human-Centric Video Annotation Pipeline
Creating task-specific questions for the aforementioned tasks hinges on extensive human-centric annotations within videos. Our video annotation pipeline, illustrated in Figure 2, emphasizes multi-modal, granular annotation pathways, supported by the Data-Juicer [6, 9] framework. While benefiting from existing operators, we’ve also developed innovative operators to enhance the open-source community. Detailed annotations and examples are in the Appendix Section 11.
3.2.1 Collecting Videos Containing People
We sourced copyright-free videos from Pexels [43], splitting each video by scene transitions using video_split_by_scene_mapper for accurate human tracking. After a series of attribute-based filtering such as duration and optical flow, a video_face_ratio_filter gauges the visibility of faces, retaining videos where a face appears in most frames for further subsequent annotation.
3.2.2 Video Mappers
We begin by implementing a specially designed operator named video_human_tracks_extraction_mapper to track each person across consecutive frames. The operator constructs tracks by iteratively detecting face and body bounding boxes, using overlap thresholds between consecutive frames to ensure accurate and consistent person localization. Additionally, this process enables an approximate count of the number of people appearing within the continuous shot. The generated tracks serve as a foundation for subsequent tasks, including person highlighting and detailed descriptions of their appearances or actions.
Next, visual attributes and descriptions for each individual are derived from track information. Specifically, the human_demographics_mapper extracts facial images from face bounding box tracks and employs existing facial attribute detection models to infer demographic labels (e.g., age, gender, race). Additionally, bounding box data at various positions (e.g., full-body or face-only) enables flexible cropping to generate individual-focused videos. This allows the description model to concentrate on appearance and posture (via the video_human_description_mapper) or facial expressions and transformations (via the video_facial_description_mapper). These processes effectively prevent the description model from misidentifying bounding box contexts or being disrupted by other individuals in the frame, while ensuring facial details are preserved during resizing.
Further, audio operations enrich video content by analyzing people-related information. The annotation process initiates with the audio_tagging_mapper classifying sound types. For speech-detected videos, specialized operators are activated: the active_speaker_detection_mapper identifies active speakers by fusing audio and visual cues, the asr_mapper transcribes speech content, the speech_emotion_recognition_mapper extracts emotional cues, and the voice_demographics_mapper profiles voice characteristics (e.g., gender, age).
Finally, annotations are extended to capture event atmosphere and broader video narratives through the video_description_mapper. These detailed annotations facilitate versatile, automated question synthesis. For example, in the Audio-Visual Speaker Matching task construction, videos are filtered by verifying that exactly one individual’s Label-3 (visual demographics) matches the video’s Label-10 (audio demographics), with all other individuals exhibiting mismatches. The matching individual is established as the ground truth, with others marked as interference, while Label-1 tracking data assigns distinct identifiers to each person. Complete task construction details are provided in the Appendix Section 12.

### v3

# HumanVBench: Probing Human-Centric Video Understanding in MLLMs with Automatically Synthesized Benchmarks

@@ Abstract
Evaluating the nuanced human-centric video understanding capabilities of Multimodal Large Language Models (MLLMs) remains a great challenge, as existing benchmarks often overlook the intricacies of emotion, behavior, and cross-modal alignment. We introduce HumanVBench, a comprehensive video benchmark designed to rigorously probe these capabilities across 16 fine-grained tasks. A cornerstone of our work is a novel and scalable benchmark construction methodology, featuring two automated pipelines that synthesize high-quality video annotations and challenging multiple-choice questions with minimal human labor. By leveraging state-of-the-art models for annotation and systematically converting model-induced errors into plausible distractors, our framework provides a generalizable “machine” for creating nuanced evaluation suites. Our extensive evaluation of 30 leading MLLMs on HumanVBench reveals critical deficiencies, particularly in perceiving subtle emotions and aligning speech with visual cues, with even top proprietary models falling short of human performance. We open-source HumanVBench and our synthesis pipelines to catalyze the development of more socially intelligent and capable video MLLMs.
Figure 1: HumanVBench is a benchmark for human-centric video understanding that features 16 fine-grained tasks, each denoted by its acronym and the number of included QA instances (middle blue box). It is synthesized with minimal human intervention by using a Video Annotation Pipeline (upper left, purple box) and a Distractor-Included QA Synthesis Pipeline (lower left, green box).
Corresponding author.
@@ 1 Introduction
In recent years, Multimodal Large Language Models have emerged as pivotal advancements, extending traditional language models to process diverse modalities such as text, images, and videos [40, 27, 22, 23, 26, 24]. Video-oriented MLLMs, in particular, have drawn growing interest for their potential to interpret video content in ways closely aligned with human perception [46]. However, the extent to which these models truly achieve human-like understanding, especially in complex human-centric scenarios, remains an open question.
Human-centric scenes in videos are rich in emotional, behavioral, and interactive content, requiring robust foundational human perceptual skills from video understanding models. Existing MLLM benchmarks mainly target general content comprehension, but often overlooking the fine-grained and multi-dimensional evaluation of essential human perceptual abilities, like nuanced understanding of emotions, behaviors, and person recognition. Synchronizing speech with visual cues, a key aspect of human perception, is also frequently neglected. Unlike humans who easily detect audio-visual mismatches, models struggle with speaker identification and lip-sync alignment. These foundational skills underpin advanced human-related reasoning tasks, including narrative reasoning, intention inference, and social intelligence. Therefore, a foundational human-centric video understanding benchmark is necessary, which requires not only new evaluation dataset but also scalable and systematic frameworks for its creation.
To address this, we introduce a novel and scalable framework for automated video benchmark creation. This framework is powered by two core pipelines: 1) the Human-Centric Video Annotation Pipeline, which leverages over twenty state-of-the-art (SOTA) data processing operators to produce dense, multimodal, automated annotations; and 2) the Distractor-Included QA Synthesis Pipeline, which generates high-quality multiple-choice questions with semantically plausible distractors. As a powerful instantiation of our framework, we present HumanVBench, a pioneering benchmark featuring 16 fine-grained QA tasks specifically tailored for human-centric analysis in video MLLMs. These tasks are organized around two core dimensions: inner emotion and outer manifestation, as depicted in Figure 1.
A key innovation of our framework lies in its model-driven design. Unlike conventional benchmarks that rely on laborious human annotation, our approach automates the multi-modal, fine-grained annotation process by integrating diverse task-specific algorithms and models. The Distractor-Included QA Synthesis Pipeline further ensures the challenging nature of the questions by leveraging multi-model ensembles and their common failure modes to generate semantically deceptive distractors. The strategy transforms the human role from manual creation to efficient quality review, significantly boosting both construction efficiency and scalability. Moreover, our method is easily adaptable to “in-the-wild” video data, enabling benchmarks creation beyond controlled or domain-specific settings.
@@
Through HumanVBench, we comprehensively evaluate 30 state-of-the-art video MLLMs, covering both open-source and commercial models. Our results reveal notable gaps between current model capabilities and human-level video understanding, particularly in nuanced emotion perception. While the proprietary models demonstrate closer human-like accuracy, open-source models often misclassify emotions due to temporal noise, underscoring the need for further architectural improvements and refined datasets.
@@
• We introduce HumanVBench, a novel video benchmark for MLLMs that emphasizes fine-grained human comprehension in videos, focusing on emotion perception, person recognition, behavioral analysis, and speech-visual alignment.
• We propose a novel and scalable framework for automated video benchmark construction, featuring two advanced pipelines for multi-modal annotation and distractor-aware QA synthesis. Our framework is generalizable and significantly reduces manual annotation effort.
• Our comprehensive evaluation of 30 SOTA video MLLMs offers key insights, facilitating in-depth discussions regarding their performance, strengths, and areas for enhancement.
• We release our benchmark, including data, evaluation, and synthesis codes at https://github.com/datajuicer/data-juicer/tree/HumanVBench , to foster further evolution of future human-centric video analysis systems.
@@ 2 Related Works
Video Benchmarks for MLLMs. Recent years have witnessed the rapid development of benchmarks for evaluating MLLMs, particularly in the video domain. Representative efforts [34, 17, 60, 65] such as Video-MME primarily target general video understanding abilities like temporal reasoning and broad scene comprehension. Although these include some human-centric content, their evaluation is often dispersed and lacks the structured, fine-grained focus necessary for deep diagnostics of human-centric perception.
To better assess human understanding, benchmarks often focus on observable activities like action recognition, pose estimation, or motion analysis [53, 45, 20], but typically neglect affective states, person recognition, or multimodal synchronization. In video emotion understanding, emotion-centric datasets [4, 39, 42] like VEATIC mostly rely on discrete emotion classification with fixed emotion categories, lacking multi-dimensional tasks such as emotion dynamics or cross-person intensity comparison. Broader MLLM benchmarks with emotion tasks often lack clear granularity and rigorous evaluation protocols. At a higher cognitive level, benchmarks [29, 30, 56] like Social-IQ target narrative reasoning, intention inference, and social intelligence, intertwining perception with higher-order reasoning and thus not isolating foundational perceptual competence.
In contrast, HumanVBench systematically targets the foundational perceptual layer of human-centric video understanding. It offers a comprehensive suite of 16 fine-grained tasks spanning person recognition, behavior and emotion perception, and audio-visual alignment, covering around 2.5k carefully synthesized multi-choice questions. This structured and multimodal evaluation serves as a prerequisite for higher-level cognitive reasoning. Quantitative comparisons with existing benchmarks are in Appendix B.
Benchmark Construction and QA Synthesis. Methods for constructing QA pairs in existing video benchmarks fall into three categories: Some use fully manual annotation, with human-written questions and answers [51, 17, 60]. Others repurpose existing datasets, as in ActivityNet-QA [54] built from ActivityNet [11], or HumanBench [45], which aggregates multiple public datasets for tasks like pose estimation and attribute recognition. Some benchmarks also incorporate automated QA generation, either by using commercial LLMs to produce QA pairs directly [32], or by synthesizing reasoning and temporal questions from existing annotations like captions, temporal labels, or scene graphs [34, 55, 44]. However, these approaches typically rely on extensive manual or historical annotations and struggle to scale to in-the-wild video data from scratch. Moreover, the automated generation of challenging distractors remains underexplored. Existing methods—often based on general semantic similarity [43, 28, 41, 16] or a single fine-tuned LLM for distractor generation [38, 5], lack mechanisms to systematically capture models’ actual inference weaknesses in the complex video domain.
To address these limitations, our method introduces a scalable, model-driven framework that constructs benchmarks directly from raw videos. This framework leverages a multi-operator approach for dense, fine-grained annotation and employs an MLLM ensemble to synthesize challenging and discriminative evaluations.
@@ 3.1 Task Design and Definition
Human observers naturally focus on individuals in videos, attending to their appearance, emotions, and behaviors. This intrinsic focus underpins our design of 16 fine-grained, human-centric tasks, aimed at evaluating MLLMs’ ability to mimic human-like perception in video analysis (Figure 1). Each task definitions and examples are provided in Appendix J. The tasks can be grouped into two categories based on content observability: Inner Emotion and Outer Manifestation.
Inner emotions are often less observable in real-world videos, requiring abstract skills to interpret subtle facial expressions, body language, and verbal cues. The tasks in this category, which we term Emotion Perception, evaluate MLLMs’ ability to detect and interpret these emotional nuances. Specifically, this category includes: Emotion Recognition (ER) identifies the most fitting emotional description for an individual. Emotion Temporal Analysis (ETA) tracks emotional changes of a person over time. Attitude Recognition (AT) assesses the inferred attitude (positive, negative, or neutral) of individuals toward specific events or entities. Emotion Intensity Comparison (EIC) differentiates and quantifies the emotional intensity of various individuals.
In contrast to inner emotions, outer manifestations deal with more tangible aspects such as identifying individual(s), causality reasoning, and synchronization of video elements like speech or singing. Guided by these aspects of human observation, we formulated three task categories:
Person Recognition evaluates the model’s ability to identify individuals in complex scenes, similar to “person-finding”: Text-to-Human (T2H) identifies a person based on a textual description. Human-to-Text (H2T) assesses the accuracy of a textual explanation attributed to a target person. Human Counting (HC) detects, tracks, and counts distinct individuals. Appearance Time Detection (ATD) identifies the specific time a specified individual appears.
Behavior Analysis scrutinizes a model’s ability to understand and analyze individual behaviors, with four tasks: Behavior Temporal Analysis (BTA) tracks changes in a specified individual’s behaviors over time. Behavior Causality Analysis (BCA) examines causal reasoning of behaviors in the video context. Action at Specified Time (AST) identifies the precise actions occurring at a given time. Time of Specific Action (TSA) detects the exact moments when specific actions occur.
Speech-Visual Alignment is a critical cross-modal category designed to test capabilities like speaker identification and lip-syncing. This category includes: Audio-Visual Speaker Matching (AVSM) correlates audio features to identify individuals and their visual appearance (gender, age). Active Speaker Detection (ASD) identifies the individual currently speaking by integrating visual and audio inputs. Audio-Visual Alignment Detection (AVAD) detects synchronization points and coherence between lip movements and audio. Speech Content Matching (SCM) analyses spoken content against text, evaluating transcription or lip-reading capabilities.
Together, these tasks offer a comprehensive assessment of MLLMs’ emotion, identity, behavior, and speech-visual alignment understanding, driving towards human-like video comprehension.
@@
Figure 3: The Distractor-Included QA Synthesis Pipeline facilitates four steps: selecting “question videos”, generating Questions, ensemble-based options generation with model-derived errors as distractors, and manually verifying multiple-choice questions.
@@ 3.2 Human-Centric Video Annotation Pipeline
Creating task-specific questions for the aforementioned tasks hinges on extensive human-centric annotations within videos. Our video annotation pipeline, illustrated in Figure 2, emphasizes multi-modal, granular annotation pathways, supported by the Data-Juicer framework [6, 8]. While benefiting from existing operators, we’ve also developed innovative operators to enhance the open-source community. Detailed annotations and examples are in the Appendix K.
Collecting Videos Containing People. We primarily sourced copyright-free videos from Pexels, which offers diverse in-the-wild footage under varied scenes and capture conditions. Movies from the MF2 benchmark [57], licensed under Public Domain 1.0, were also included to further increase content diversity. Importantly, we used only the raw visual data without any pre-existing captions or metadata. All annotations were generated internally by our pipelines, thereby preventing risks of data leakage. We begin by splitting each video into scenes with video_split_by_scene_mapper to facilitate subsequent accurate human tracking. The resulting clips are then filtered by attributes such as duration and optical flow, after which video_face_ratio_filter is applied to ensure that faces are visible in most frames. This guarantees that the selected videos are primarily human-centric and ready for further annotation.
Human-Centric Video Annotation Pipeline. We design a set of operators to extract and enrich human-centered video information. First, video_human_tracks_extraction_mapper links detected faces and bodies across consecutive frames using overlap thresholds, producing reliable person tracks and an approximate count of individuals in a shot. These tracks support subsequent tasks such as person highlighting and appearance/action descriptions.
From the tracks, human_demographics_mapper extracts face crops and applies attribute models to infer demographic labels (e.g., age, gender, race). Flexible bounding box crops (full-body or face-only) further allow generating individual-focused clips. Based on this, video_human_description_mapper captures appearance and posture using an MLLM, while video_facial_description_mapper focuses on facial expressions and changes. This ensures descriptions are unaffected by background or other people, while preserving facial detail.
For audio, audio_tagging_mapper first classifies sound types. If speech is present, additional operators are invoked: active_speaker_detection_mapper fuses audio-visual cues to localize speakers, asr_mapper transcribes speech, speech_emotion_recognition_mapper detects emotions, and voice_demographics_mapper profiles vocal attributes (e.g., gender, age).
Finally, video_description_mapper summarizes overall atmosphere and narrative using MLLMs. These multimodal annotations enable automated task construction. For example, in the Audio-Visual Speaker Matching task, videos are selected where exactly one individual’s visual demographics (Label-3) match the audio demographics (Label-10), while other individuals’ features mismatch. The matched individual serves as ground truth, with IDs assigned via Label-1 tracking. Implementation details for each operator and full task construction appear in Appendices K and L.